OpenStudy (anonymous):

@TheSmartOne Can you help me fill in the formula for this one too? What is the length of the segment joining the points at (–6, 8) and (6, 3)? _________ units

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

Sure thing. One sec :)

TheSmartOne (thesmartone):

\(\large\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\) \(\sf x_1 = -6 \\ x_2 =6 \\y_1=8 \\y_2=3\) \(\large\sqrt{(6-(-6))^2+(3-8)^2}\)

OpenStudy (anonymous):

13?

OpenStudy (anonymous):

@Destinymasha @sammixboo @TheEdwardsFamily @pooja195 @dan815

OpenStudy (anonymous):

yes 13